Multiplication of two numbers that differ by 6 (This trick only works if you have memorized or can quickly calculate the squares of numbers.  If the two numbers differ by 6 then their product is the square of their average minus 9.


Let me explain this rule by taking examples


10*16 = 13^2 - 9 = 160


22*28 = 25^2 - 9 = 616


Grade:12

1 Answers

Nishant Vora IIT Patna
9 years ago
The Explanation is simple

10*16 = (13-3)*(13+3) = 13^2 - 3^2 = 13^2 - 9
Hence Proved
_________________________________________________________________________________________________________________________________________________________________________________________________________________________________________